write an equation of the line parallel to the line y=3/4x-6 and having to y intercept of (0,-2)
Answer by ewatrrr(24785)


Hi
Using the standard slope-intercept form for an equation of a line y = mx + b
where m is the slope and b the y-intercept.
line parallel to the line y = 3/4x-6 m = (3/4)
parallel Lines have identical slopes
y = (3/4)x -2 | Stated y-intercept of (0,-2)